The difference of two real numbers is 7, If the difference of their squares is 210. then the sum of those two real numbers is
Answer by stanbon(75887)

The difference of two real numbers is 7, If the difference of their squares is 210. then the sum of those two real numbers is
------------------
Equations:
x^2 - y^2 = 210
x - y = 7
-----
Factor the 1st to get:
(x-y)(x+y) = 210
-----
Substitute for (x-y) to get:
7(x+y) = 210
Then x + y = 30
And x - y = 7
-----
Add and solve for "x":
2x = 37
x = 18.5
----
Solve for "y"::
x + y = 30
18.5 + y = 30
y = 11.5
-------------
Cheers,
Stan H.
--------------